# Efficient and Timely Memory Access

### v1

@@ I Introduction
This work examines status updating systems in which sources generate time-stamped status updates of a process of interest, and these updates are stored/written in a memory system. A reader fulfills clients’ requests for these updates by reading from the memory. The asynchronous nature of reader-writer interactions within memory systems introduces significant challenges. In particular, the readers’ memory accesses should be optimized for timely processing of source updates as the reader becomes aware of fresher updates in the memory only when it chooses to query the memory. Furthermore, the memory access process must be regulated by a synchronization method between readers and writers to avoid race conditions.
@@ I-A Related Work
In [3], the authors addressed the timely processing of updates stored in memory from multiple sources. The framework involved a system where the Reader samples the memory as a renewal point process. Under this model, it was shown that a lazy sampling policy from the Reader proved to be an optimal strategy. The rationale behind this finding was that lazy sampling effectively mitigates the negative impacts of high variance in the client’s processing times. This work differs from [3] in that it no longer assumes renewal sampling from the Reader and in contrast to previous work, in this study, the Reader pays a cost of sampling. Nevertheless, the primary objective remains consistent: to identify an optimal reading policy for the Reader that maximizes the timeliness of source updates with the client.
We note that the concept of timely memory sampling, wherein the Reader incurs a cost for sampling for age reduction, shares similarities with research focused on managing access for multiple users within a communication channel. Various studies in the AoI literature have explored Whittle’s index-based transmission scheduling algorithms [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], wherein the scheduling problem is decomposed into multiple independent subproblems. Within each subproblem, an additional cost $(C)$ is associated with updating the user. The goal is to determine, in each time slot, whether updating the user is warranted, thereby striking a balance between the updating cost and the age-related costs.
@@ II System Model
We consider a discrete-time slotted system with slots labelled $t=0,1,2,\ldots$. The system involves two key processes: writing the time-varying data from the source into the memory and reading the source data from memory. The modeling details of these processes are discussed below.
@@ II-A Writing source updates to memory
II-B Sampling source updates from memory
@@
The Reader generates an age process $y(t)$ at the input to the client that is a sampled version of source update age process $x(t)$ in the memory. Hence we say the Reader samples the updates in the memory.
@@ II-C Markov Decision Process Formulation
  • Action: Let $a(t)\in\{0,1\}$ denote the action taken in slot $t$ indicating Reader’s decision, where $a(t)=1$ if Reader decides to read and $a(t)=0$ if idle.
@@
  • Cost: The cost $C(s(t);a(t))$ incurred in state $s(t)$ in time slot $t$ under action $a(t)$ is defined as:
    $C(s(t)=(x,y);~{}a(t)=a)\coloneqq y+ca.$ (2)
Let $\pi=\{a(0),a(1),\ldots\}$ denote a policy that specifies an action $a(t)$ at slot $t$. The expected average cost under policy $\pi$ starting from a given initial state at $t=0$, $s(0)=s$, is defined as:
$g_{\pi}(s)$ $=\limsup_{{T\to\infty}}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}C(s(t);a(t))\mid s(0)=s\right],$
$=\limsup_{{T\to\infty}}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}(y(t)+ca(t))\right].$ (3)
@@
Our cost minimization problem falls within the category of average cost minimization problems. Given that the age can grow unbounded, both the number of states and the cost in each stage are countably infinite. In such Markov Decision Processes (MDPs), the existence of an optimal policy, whether stationary or non-stationary, is not guaranteed [17, Chap 5]. Notably, even the existence of an optimal stationary policy may not hold, while an optimal non-stationary policy might exist [18].
Analyzing average cost problems with an infinite state space poses inherent difficulties. However, under certain conditions and structures, it is possible to develop useful results. Proving the existence of an optimal average cost stationary policy is not an immediate goal in this paper and we defer this discussion to later in Section V. There, we draw upon results from [19], which provides conditions ensuring the existence of an expected average cost optimal stationary policy. We verify that these conditions hold for our problem. In the subsequent section, we derive results regarding the structure of the optimal policy under the assumption that the optimal policy exists and the relative cost Bellman’s equation is valid.
@@ III-A Discounted Cost
The proof, using mathematical induction on (5), is straightforward but omitted because of space constraints.
@@
Proof of this proposition is provided in the Appendix. Another version of this proposition asserts that if the $\alpha$-optimal action is to sample in state $(x,y)$ at stage $n$, then it is also optimal to sample in every $(x,y^{\prime})$ with $y^{\prime}\geq y$ at stage $n$. The proof employing the value iteration (5) is omitted as it is similar to that of Proposition 2.
@@
The proof appears in the Appendix. The intuitive structure of the optimal policy is that with knowledge of the age in the memory, the Reader should refrain from sampling if the reduction in age doesn’t justify the sampling cost. To further characterize this intuition, we introduce the following proposition. The proof appears in the Appendix.
@@
Specifically, when the memory is freshly updated, the Reader must assess whether sampling is worthwhile. If it opts against sampling initially, it should consistently abstain from sampling in subsequent slots until the memory undergoes another update, as the age reduction remains constant in the absence of changes. In terms of the MDP $\mathcal{M}$, this concept translates to making a decision in the state $(0,y)$. If the optimal decision is not to sample at this point, then the Reader should consistently refrain from sampling in states $(1,y+1)$, $(2,y+2)$, and so on.
@@ III-B Average Cost Optimality
Since the conditions of Theorem 2 (in section V) hold, the cost-optimal policy $\pi^{*}$ is the limit point of $\alpha$-optimal policies $\pi^{*}_{\alpha}$ with $\alpha\to 1$ [19, Lemma]. Therefore, Propositions 2 and 4 are sufficient to provide the structure of average cost optimal policy. Specifically, Propositions 2 and 4 imply that there exists a threshold $Y_{0}$ such that it is optimal to sample in $(0,y)$ for every $y\geq Y_{0}$ and idle otherwise.
At this point, it is important to mention the set of feasible states under $\pi^{*}$. With $Y_{0}=1$, the optimal policy dictates sampling in every state $(0,y)$ where $y\geq 1$. Upon sampling in $(0,1)$, the system transitions to feasible states, specifically $\{(0,1),(1,1)\}$. In state $(1,1)$, a close examination of Bellman’s equation (4) reveals that it is optimal to idle. Therefore, the set of possible states when choosing to idle in $(1,1)$ becomes $\{(0,2),(2,2)\}$. Subsequent transitions follow a pattern where sampling in $(0,y)$ leads to states $\{(0,1),(1,1)\}$, and choosing to idle in states $(i,i)$ with $i\in\mathbb{N}$ resulting in $\{(0,i+1),(i+1,i+1)\}$.
@@
The proof appears in the Appendix. We now use Proposition 6 to derive the optimal threshold.
@@
Figure 2: Plot of average cost $g_{0}(Y_{0})$ as a function of threshold $Y_{0}$ with sampling cost $c=80$. The circle point is the minimum of $g_{0}(Y_{0})$, the cross is the approximate optimal threshold $\tilde{Y}_{0}^{*}$.
@@ IV Numerical Evaluation
Figure 2 illustrates the behavior of the average cost given by (10) with respect to $Y_{0}$. For fixed system parameters $p$ and $c$, the plot reveals an initial decline in average cost as threshold $Y_{0}$ increases. This trend aligns with expectations, as a low threshold prompts the Reader to sample too frequently, incurring the sampling cost without achieving a significant reduction in age, ultimately leading to a higher average cost. As $Y_{0}$ increases further, the cost of sampling approaches the gain obtained with age reduction. However, setting threshold $Y_{0}$ too high precludes timely access to the memory, resulting in increased age at the client and consequently increased average cost of the system. The observed behavior in Figure 2 highlights the existence of an optimal threshold at which the cost payed for sampling is justified with the corresponding reduction in age.
@@
Fig. 3 illustrates the value of optimal threshold $Y_{0}^{*}$ as a function of probability $p$ of source update publication in a slot. We observe that the optimal threshold increases with $p$. When the Reader is required to make a decision in a given slot, it assesses both the age at the client and the age in the memory. These evaluations contribute to determining the potential age reduction vs the cost of sampling. In scenarios where the client’s update is deemed sufficiently recent, the Reader may choose to skip sampling. This decision is influenced by a higher probability ($p$) of obtaining a more recent update soon, that will perhaps be worth sampling.
Figure 4 compares the optimal average cost $g$ with the lower bound provided in (13). The tightness of the lower bound is evident, as it closely aligns with the curve of the optimal average cost. Additionally, the figure illustrates that the optimal cost tends to increase with an increase in the sampling cost $c$. This suggests that while designing the cost structure, the cost should be sufficiently high but not excessively so. Furthermore, the plot shows that the average cost decreases as the probability $p$ of memory updates in a slot increases. This is intuitive, as frequent memory updates increase the likelihood of the Reader receiving a fresh update when it samples, thereby reducing the age at the client.
@@ V Stationary average cost optimal policy
In this section we verify that the average cost optimality equation for MDP holds for $\mathcal{M}$. To get started, we need the following result.
@@
Under the deterministic stationary policy $\theta$ of reading in every slot, the system exhibits an irreducible, ergodic Markov Chain, with expected cost $M(x,y)$ of first passage from state $s=(x,y)$ to $(0,1)$ satisfying
@@
Proof of Lemma 3 appears in the Appendix. We now employ the lemma in verifying the conditions of the following theorem.
@@ VI Conclusion
This paper focused on a class of systems where source updates are disseminated using shared memory. The Writer process records these source updates in the memory, and a Reader fulfills clients’ requests for these measurements by reading from the memory. We studied the problem of optimizing memory access by the Reader with respect to minimizing average cost. Our main contributions included establishing the existence of an optimal stationary deterministic policy for our Markov Decision Process (MDP). Furthermore, we demonstrated that the optimal policy has a threshold structure.
A key insight from our analysis was that the Reader should choose to sample only when the memory undergoes an update. If the Reader decides not to sample, this decision of staying idle should perpetuate in subsequent slots until the memory is updated with a fresh source update. This is because, in the absence of updates, there is no change in age reduction; it remains the same as when the memory was last updated. Finally, an interesting extension to this work would involve investigating an alternate regime for optimizing memory access, where the Reader operates without knowledge of source update age in the memory.

### v2

@@ I Introduction
This work examines status updating systems in which sources generate time-stamped status updates of a process of interest, and these updates are stored/written in a memory system. A reader fulfills clients’ requests for these updates by reading/sampling from the memory. The asynchronous nature of reader-writer interactions within memory systems introduces significant challenges. In particular, the readers’ memory accesses should be optimized for timely processing of source updates as the reader becomes aware of fresher updates in the memory only when it chooses to query the memory. Furthermore, the memory access process must be regulated by a synchronization method between readers and writers to avoid race conditions.
@@ I-A Related Work
In [3], the authors addressed the timely processing of updates stored in memory from multiple sources. They proposed a system where the reader samples memory as a renewal point process, showing that a lazy sampling policy was optimal. The rationale was that lazy sampling mitigated negative impacts of high variance in client processing times. This work differs from [3] as it no longer assumes renewal sampling and introduces a cost for sampling. Nevertheless, the primary objective remains consistent: to identify an optimal reading policy maximizing source update timeliness with the client.
We note that the concept of timely memory sampling, wherein the reader incurs a cost for sampling for age reduction, shares similarities with research focused on managing access for multiple users within a communication channel. Various studies in the AoI literature have explored Whittle’s index-based transmission scheduling algorithms [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], where the scheduling problem is decomposed into multiple independent subproblems. Within each subproblem, an additional cost $(C)$ is associated with updating the user.
However, there is a conceptual difference in the cost associated with the decoupled problem and this study. In a Whittle index policy, the minimum cost that makes both actions — updating a user or idling — equally desirable is used as a mechanism to choose one of the many users. In this work, we enforce an explicit cost of accessing the memory, and we study the trade-offs observed with age and memory access by varying this system parameter. However, it is not a mechanism designed to distinguish between users.
@@ II System Model
We consider a discrete-time slotted system with slots labelled $t=0,1,2,\ldots$. The modeling details of writing and reading processes are discussed below.
@@ II-A Writing source updates to memory
II-B Reading source updates from memory
@@
The Reader generates an age process $y(t)$ at the input to the client that is a sampled version of source update age process $x(t)$ in the memory. Hence, we say the Reader samples the updates in the memory.
@@ II-C Markov Decision Process Formulation
  • Action: Let $a(t)\in A=\{0,1\}$ denote the action taken in slot $t$ indicating Reader’s decision, where $a(t)=1$ if Reader decides to read and $a(t)=0$ if idle.
@@
  • Cost: The cost $C(s(t);a(t))$ incurred in state $s(t)$ in time slot $t$ under action $a(t)$ is defined as $C(s(t)=(x,y);~{}a(t)=a)\coloneqq y+ca$.
Let $\pi:S\rightarrow A$ denote a policy that for each state $s(t)\in S$ specifies an action $a(t)=\pi(s(t))\in A$ at slot $t$. The expected average cost under policy $\pi$ starting from a given initial state at $t=0$, $s(0)=(x,y)$, is defined as:
$g_{\pi}(x,y)=\limsup_{{T\to\infty}}\frac{1}{T}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}(y(t)+ca(t))\right]$ (2)
@@
Our cost minimization problem falls within the category of average cost minimization problems. Given that the age can grow unbounded, both the number of states and the cost in each stage are countably infinite. In such MDPs, the existence of an optimal policy, whether stationary or non-stationary, is not guaranteed [17, Chap 5]. Notably, even the existence of an optimal stationary policy may not hold, while an optimal non-stationary policy might exist [18].
Proving the existence of an optimal average cost stationary policy is not an immediate goal in this paper and we defer this discussion to later in Section V. There, we draw upon results from [19], which provides conditions ensuring the existence of an expected average cost optimal stationary policy. We verify that these conditions hold for our problem. In the subsequent section, we derive results regarding the structure of the optimal policy under the assumption that the optimal policy exists and the relative cost Bellman’s equation is valid.
@@ III-A Discounted Cost
The proof, which uses mathematical induction on (4), is straightforward and has been omitted. From here, proofs, if not shown, appear in the Appendix.
@@
Another version of this proposition asserts that if the $\alpha$-optimal action is to sample in state $(x,y)$ at stage $n$, then it is also optimal to sample in every $(x,y^{\prime})$ with $y^{\prime}\geq y$ at stage $n$. The proof employing the value iteration (4) is omitted as it is similar to that of Proposition 2.
@@
The intuitive structure of the optimal policy is that with knowledge of the age in the memory, the Reader should refrain from sampling if the reduction in age doesn’t justify the sampling cost. To further characterize this intuition, we introduce the following proposition.
@@
Specifically, when the memory is freshly updated, the Reader must assess whether sampling is worthwhile. If it opts against sampling initially, it should consistently abstain from sampling in subsequent slots until the memory undergoes another update, as the age reduction remains constant in the absence of changes. In terms of the MDP $\mathcal{M}$, this concept translates to making a decision in the state $(0,y)$. If the optimal decision is not to sample in $(0,y)$, then the Reader should consistently refrain from sampling in states $(1,y+1)$, $(2,y+2)$, and so on.
@@ III-B Average Cost Optimality
Since the conditions of Theorem 2 (in section V) hold, the average-cost optimal policy $\pi^{*}$ is the limit point of $\alpha$-optimal policies $\pi^{*}_{\alpha}$ with $\alpha\to 1$ [19, Lemma]. Therefore, Propositions 2 and 4 are sufficient to provide the structure of average cost optimal policy. Specifically, Propositions 2 and 4 imply that there exists a threshold $Y_{0}$ such that it is optimal to sample in $(0,y)$ for every $y\geq Y_{0}$ and idle otherwise.
At this point, it is important to mention the set of feasible states under $\pi^{*}$. With $Y_{0}=1$, the optimal policy dictates sampling in every state $(0,y)$ with $y\geq 1$. Upon sampling in $(0,1)$, the system transitions to feasible states, specifically $\{(0,1),(1,1)\}$. In state $(1,1)$, a close examination of Bellman’s equation (3) reveals that it is optimal to idle. Therefore, the set of possible states when choosing to idle in $(1,1)$ becomes $\{(0,2),(2,2)\}$. Subsequent transitions follow a pattern where sampling in $(0,y)$ leads to states $\{(0,1),(1,1)\}$, and choosing to idle in states $(i,i)$ with $i\in\mathbb{N}$ resulting in $\{(0,i+1),(i+1,i+1)\}$.
@@
We use Proposition 6 to derive the optimal threshold.
@@
Theorem 1 provides an explicit expression for the optimal threshold $Y_{0}^{*}$. However, evaluating the optimal average cost $g=g_{0}(Y_{0}^{*})$ using (9) doesn’t directly show the relationship between system parameters $c$ and $p$. The following lemma provides a close approximation to the optimal average cost and captures the impact of these key system parameters on the average cost.
@@
Figure 2: Plot of average cost $g_{0}(Y_{0})$ as a function of threshold $Y_{0}$ with sampling cost $c=80$. Here, $\circ$ is the true optimal cost $g_{0}(Y_{0}^{*})$, and $\times$ is the approximate optimal average cost $g_{0}(\tilde{Y}_{0}^{*})$.
@@ IV Numerical Evaluation
Fig. 2 shows how the average cost $g_{0}(Y_{0})$, given by (9), changes with threshold $Y_{0}$. Initially, as $Y_{0}$ increases, the average cost decreases. This is because a low threshold leads to excessive sampling, incurring costs without much age reduction, resulting in a higher average cost. As $Y_{0}$ increases further, the cost of sampling approaches the gain in age reduction. However, setting $Y_{0}$ too high delays memory access, increasing client age and consequently the average cost. Fig. 2 highlights the existence of an optimal threshold where the cost of sampling justifies the age reduction.
Fig. 3 illustrates the optimal threshold $Y_{0}^{*}$ as a function of the source update probability $p$. We observe that the optimal threshold increases with $p$. When the Reader is required to make a decision in a given slot, it assesses both the age at the client and the age in the memory. These evaluations contribute to determining the potential age reduction vs the cost of sampling. In scenarios where the client’s update is deemed sufficiently recent, the Reader may choose to skip sampling. This decision is influenced by a higher probability ($p$) of obtaining a more recent update soon, that will perhaps be worth sampling.
Fig. 4 compares the optimal average cost $g$ with the lower bound from (12). The tightness of the lower bound is evident, as it closely aligns with the curve of the optimal average cost. The plot also demonstrates that $g$ increases with sampling costs $c$. This suggests that while designing the cost structure, the cost should be sufficiently high but not excessively so. Moreover, $g$ decreases as the update probability $p$ increases. This is intuitive, as frequent memory updates increase the likelihood of the Reader receiving a fresh update when it samples, thereby reducing the age at the client.
@@ V Stationary average cost optimal policy
In this section, we verify that the average cost optimality equation for MDP holds for $\mathcal{M}$. To get started, we need the following result.
@@
Under the deterministic stationary policy of reading in every slot, the system exhibits an irreducible, ergodic Markov Chain, with expected cost $M(x,y)$ of first passage from state $s=(x,y)$ to $(0,1)$ satisfying
@@
We now employ Lemma 3 in verifying the conditions of the following theorem.
@@ VI Conclusion
This paper focused on a class of systems where source updates are disseminated using shared memory. The Writer process records these source updates in the memory, and a Reader fulfills clients’ requests for these measurements by reading from the memory. We addressed optimizing memory access to minimize average cost, establishing the existence of an optimal stationary deterministic policy for our Markov Decision Process (MDP) and showing its threshold structure.
We found that in the optimal policy, if the Reader decides to stay idle when the memory undergoes an update, then the optimal action is to not sample in subsequent slots until the memory is updated with a fresh source update. Finally, we observe that the age achieved by this optimal policy serves as a lower bound on the achievable average age when the Reader is unaware of the memory state. Investigating the optimal policy structure for this scenario, where the Reader does not know when the memory is updated, is left for future research.